# Code-Optimise: Self-Generated Preference Data for Correctness and Efficiency

### v1

@@ Abstract
Code Language Models have been trained to generate accurate solutions, typically with no regard for runtime. On the other hand, previous works that explored execution optimisation have observed corresponding drops in functional correctness. To that end, we introduce Code-Optimise, a framework that incorporates both correctness (passed, failed) and runtime (quick, slow) as learning signals via self-generated preference data. Our framework is both lightweight and robust as it dynamically selects solutions to reduce overfitting while avoiding a reliance on larger models for learning signals. Code-Optimise achieves significant improvements in $pass@k$ while decreasing the competitive baseline runtimes by an additional 6% for in-domain data and up to 3% for out-of-domain data. As a byproduct, the average length of the generated solutions is reduced by up to 48% on MBPP and 23% on HumanEval, resulting in faster and cheaper inference. The generated data and codebase will be open-sourced at www.open-source.link.
@@
Table 1: Statistics of our self-generated preference data. 1) A Model generates 100 solutions per problem out of Total problems in each Split. 2) Functional correctness and runtime are annotated. 3) Problems are filtered to retain those with at least 2 passing and 1 failing solution (Filtered). A low coefficient of variation (CoV $\leq$ 0.1) across 5 runs indicates that runtime measurements are stable. Ratio is the percentage of $\frac{\textbf{Filtered}}{\textbf{Total}}$ retained code solutions.
@@ 1 Introduction
Code Language Models (CLMs) trained on large code repositories such as The Stack Kocetkov et al. (2022); Lozhkov et al. (2024) gradually increase their understanding of code semantics. CLMs are thus able to generate functionally correct and reasonably efficient solutions to programming problems Austin et al. (2021); Chen et al. (2021), among many other code related skills Li et al. (2023). Shypula et al. (2023) have shown that CLMs can optimise slow-running code to achieve large runtime gains but at a substantial cost to correctness (down by up to $\sim$30%). Subsequent research has focused mostly on improving code correctness. On the data perspective, a common way of improving functional correctness is via distilled supervised fine-tuning Tunstall et al. (2023); Xu et al. (2023); Luo et al. (2023); Wei et al. (2023) on training data generated by large models such as GPT-4 Achiam et al. (2023). However, in many cases, due to legal, financial and/or privacy constraints, it is not feasible to rely on proprietary data. Additionally, we seek to overcome the limitations of supervised fine-tuning (SFT), which only optimises for ‘positive’ examples, with no means of reducing the likelihood of generating undesirable (e.g. incorrect or slow) code. Although such issues may be addressed via Reinforcement Learning (RL) Le et al. (2022); Wang et al. (2022); Gorinski et al. (2023), they often come with added complexity and instability. Therefore, we opt for Direct Preference Optimisation Rafailov et al. (2024) as our preferred fine-tuning method due to its simplicity and widespread adoption. We propose Code-Optimise, a lightweight framework that trains CLMs with our self-generated preference data for correctness (passed/failed) and efficiency (quick/slow), shown in Figure 1. Starting from a small collection of problems and unit tests, Code-Optimise bootstraps the pretrained CLM to generate the required learning signals thereby exposing the model to on-policy automatically annotated data. Code-Optimise provides additional robustness by dynamically selecting solutions during training to reduce overfitting. Our method consists of three steps: 1) Sampling; generate N solutions for each problem description, 2) Annotation; automatically label each solution for correctness and runtime, 3) Optimisation; fine-tune the CLM on the self-generated preference data using several lightweight configurations. The main contributions of Code-Optimise are:
  • We create and publish a novel code preference dataset (and a recipe to extend it) that enables multi-objective optimisation (code correctness and runtime efficiency) of CLMs.
@@
  • We demonstrate that runtimes are reduced by up to 6% for MBPP and up to 3% for HumanEval over competitive baseline CLMs. Finally, the length of generated solutions is reduced by up to 23% for HumanEval and up to 48% for MBPP, decreasing inference costs.
@@ 2 Code-Optimise
We now introduce Code-Optimise, a lightweight method for optimisation of CLMs aimed at improving functional correctness of code as well as reducing its runtime, shown in Figure 1.
@@ 2.1 Sampling
We assume access to $D_{seed}=\{x_{i},y_{i},ut_{i}\}_{i=1}^{N}$, a dataset of problem descriptions $x_{i}$ and the corresponding unit tests $ut_{i}$ that can be used for sampling and testing new solutions from the CLM, denoted $CLM_{base}$ henceforth. Since fine-tuning the model on the limited solutions $y_{i}$ would lead to overfitting, we leverage its extensive pretraining to generate a multitude of diverse solutions to obtain additional training data. We sample 100 solutions from $CLM_{base}$ for each problem description with multinomial sampling due to its lower computational cost. A temperature of $t=0.6$ is applied to achieve a balance between functional correctness and diversity, resulting in non-uniform runtimes.
@@ 2.3 Optimisation
In this step, the model is fine-tuned on $D_{train}$ to bias $CLM_{base}$ towards generating more functionally correct and runtime-efficient solutions. Although several methods for preference data optimisation exist (Yuan et al., 2023; Zhao et al., 2023; Liu et al., 2024; Azar et al., 2023; Ethayarajh et al., 2024; Hong et al., 2024), we opt for DPO due to its simplicity and wide adoption. We also benchmark SFT due to its widespread use in prior work.
@@ 3.4 Model Scaling
Figures 4, 5 and 6 show the evolution of functional correctness, runtimes and lengths of generated solutions as the number of trainable parameters increases (1B - 13B). Analysing $pass@1$ in Figure 4, we can see that larger DPO models achieve a more significant improvement over the baseline and SFT, particularly for in-domain problems. Somewhat surprisingly, functional correctness for HumanEval (out-of-domain) improves at a faster rate than MBPP (up to 7B parameters). In Figure 5, we observe that as the DPO models increase in size, their runtimes relative to the baseline remain consistent. The $DPO_{PvF}$ configuration tends to average somewhat slower runtimes as this setup only optimises for correctness thus sacrificing efficiency. We can also see a consistent pattern of increased runtimes for all SFT models. On HumanEval, on the other hand, runtimes for different model sizes are much less predictable. However, on average, our best configuration $DPO_{QvS}$ does show an improvement over the already competitive baseline CLMs.
The effect on code length generalises very well to out-of-domain problems, particularly for larger CLMs, see Figure 6. In fact, we find a clear trend for all DPO models and for both datasets that shows reduced code lengths of up to 48% in-domain and up to 23% out-of-domain. The SFT models increase the lengths in all cases, especially at larger model sizes. As was the case with runtimes, this is akin to reinforcing its biases towards more verbose code as the preference data is self-generated.
@@ 3.5 Qualitative Analysis
Figure 8 shows several solutions to a typical programming problem taken from MBPP that gives a more tangible form to our results. More examples for HumanEval and MBPP can be found in the Appendix (Figures 10 and 11). Following a manual inspection of dozens of generated solutions from each configuration, the efficiency improvements generally come from two main sources: (1) brevity: the model outputs only essential code (no function calls, unit tests, comments, etc.), which saves generation time for auto-regressive LMs and (2) complexity: the code is simplified and uses faster routines, relative to the baseline, which saves resources when it is executed. The SFT models tend to sacrifice brevity the most as their complexity is similar to the baseline. Figures 8, 10, and 11 show several examples of this, e.g. adding function calls to the newly generated solution, possibly with import statements and/or expected inputs or outputs [^1] Adding test cases to code is generally considered good programming practice. However, for the purpose of pure efficiency (our case), this can potentially add significant cost.. This is in line with the observation from Figure 3 where the SFT models appear to be more verbose and biased towards longer outputs. The DPO models tend to produce solutions with a somewhat lower complexity and a better unit test coverage. Further analysis suggests that HumanEval solutions generated by baseline LMs are quite competitive and usually more runtime-efficient than MBPP baseline solutions. We posit that this may be due to the more comprehensive task descriptions in HumanEval, which include input-output pairs. Among DPO models, we do not observe a clear winner in qualitative analysis although $DPO_{QvS}$ is the best setup in terms of aggregate results.
@@
3.6 Importance of Solution Selection
4 Related Work
Distilled Supervised Fine-Tuning
was applied to code generation models for the sole purpose of improving functional correctness. CLMs such as MagiCoder Wei et al. (2023) and WizardCoder Luo et al. (2023) query large proprietary models to provide the necessary training data Cui et al. (2023); Xu et al. (2023). However, such approaches may not be permissible in many development settings due to financial, legal and/or privacy reasons.
Reinforcement Learning
Le et al. (2022); Wang et al. (2022); Gorinski et al. (2023) can overcome the shortcomings of supervised fine-tuning by effectively propagating the negative rewards for dysfunctional code. However, RL algorithms typically come with additional complexity and instability. Recently, Rafailov et al. (2024) proposed Direct Preference Optimisation (DPO) as an alternative to Reinforcement Learning from Human Feedback for aligning language models with human preferences Tunstall et al. (2023). DPO serves as a form of offline reinforcement learning that directly optimises on a given set of trajectories without the need for a separate reward model. We should note that the aforementioned RL approaches to code synthesis only consider functional correctness and not the runtime of generated solutions.
Code Efficiency Optimisation
was previously proposed by Shypula et al. (2023) as a code editing/repair task where slow-running code had to be edited to achieve a faster runtime. Models were optimised on a newly curated dataset augmented by synthetic test cases through various methods of prompting and fine-tuning. However, the greatly reduced runtimes came at a significant cost to functional correctness. In many configurations, the model edits reduce performance by up to 30% with ‘smaller’ CLMs (7B, 13B) suffering a larger degradation. We hypothesise that this may be due to a) overfitting the single runtime objective (in contrast with our work where the aim is to optimise both correctness and runtime) and b) the removal of failed programs from the dataset leading the CLMs to struggle with semantics of correct versus incorrect code. We opt to not compare directly with this work as their method is specifically curated to the code editing task where an already functionally correct but inefficient program is assumed as input. On the contrary, we aim to produce programming solutions from scratch that are both functionally correct and runtime/inference efficient.
@@ 5 Conclusions
We have introduced Code-Optimise, a lightweight framework for improving functional correctness and runtime via self-generated code preference data optimisation (quick versus slow and passing versus failing solutions). Our experiments have shown several benefits: 1) functional correctness is significantly improved, particularly for smaller models, 2) dynamic solution selection during training provides an additional benefit by reducing overfitting, 3) runtime is reduced by up to 6% for MBPP and up to 3% for HumanEval over strong baseline CLMs, lowering the cost of code execution, 4) code length is significantly shortened, up to 48% for MBPP and up to 23% for HumanEval, which reduces inference cost and improves response times. We hope that our insights as well as our novel dataset will stimulate further exciting research in this area.
@@ 6 Limitations
Timing the execution of short programs accurately is challenging and despite our best efforts, the runtime measurements could probably be improved further with additional software engineering efforts. This would also provide a cleaner and more stable learning signal for Code-Optimise, which could potentially improve results. While our methodology is highly data-efficient, using only $\sim$200 open-source prompts for training data generation, obtaining additional high-quality problems (free from proprietary/licensing issues) may potentially yield better results. Other code-related tasks that may be amenable to optimisation for improved runtime/inference could potentially benefit from our methodology and as such may be investigated outside of the scope of this paper. While we conducted all experiments using Python, we acknowledge that less popular/similar programming languages should also be investigated in follow-up work.
@@ A.1 Dataset
The Mostly Basic Programming Problems introduced by Austin et al. (2021) consists of 974 crowd-sourced Python programming challenges. Each problem comprises a description, an example code solution and a few automated test cases. The dataset contains training, validation and test splits. We utilise the training and validation splits for optimisation, while the test split serves as the in-domain test data distribution.
@@
Chen et al. (2021) comprises 164 Python programming challenges. The function signatures, docstrings, example solutions and several unit tests were handwritten for each problem. We leverage HumanEval as our out-of-domain test set as the descriptions in MBPP do not contain any unit tests and the writing style of HumanEval problems does not follow a consistent format. This helps us evaluate robustness to handwritten prompts.
@@ A.2 Training
We use the StarCoder Li et al. (2023) and CodeLlama Rozière et al. (2024) families of models in our experiments. We opt for the pretrained (base) versions with sizes of 1B and 3B for StarCoder and 7B and 13B for CodeLlama, hosted on HuggingFace Wolf et al. (2020). During training, we fine-tune each model using a total of 30 epochs and select the best model based on the lowest validation loss. We use a learning rate of $5e^{-7}$ with a linear scheduler, a 10% warm-up, and a maximum sequence length of 2048 tokens.

### v2

@@ Abstract
Code Language Models have been trained to generate accurate solutions, typically with no regard for runtime. On the other hand, previous works that explored execution optimisation have observed corresponding drops in functional correctness. To that end, we introduce Code-Optimise, a framework that incorporates both correctness (passed, failed) and runtime (quick, slow) as learning signals via self-generated preference data. Our framework is both lightweight and robust as it dynamically selects solutions to reduce overfitting while avoiding a reliance on larger models for learning signals. Code-Optimise achieves significant improvements in $pass@k$ while decreasing the competitive baseline runtimes by an additional 6% for in-domain data and up to 3% for out-of-domain data. As a by-product, the average length of the generated solutions is reduced by up to 48% on MBPP and 23% on HumanEval, resulting in faster and cheaper inference. The generated data and codebase is open-sourced at https://github.com/huawei-noah/HEBO/tree/Code_Optimise.
@@
Table 1: Statistics of our self-generated preference data. 1) A Model generates 100 solutions per problem out of the Total problems in each Split. 2) Functional correctness and runtime are annotated. 3) Problems are filtered to retain those with at least 2 passing and 1 failing solution (Filtered). A low coefficient of variation (CoV $\leq$ 0.1) across 5 runs indicates that runtime measurements are stable. Ratio is the percentage of $\frac{\textbf{Filtered}}{\textbf{Total}}$ retained code solutions.
@@ 1 Introduction
Code Language Models (CLMs) trained on large code repositories such as The Stack Kocetkov et al. (2022); Lozhkov et al. (2024) gradually increase their understanding of code semantics. CLMs are thus able to generate functionally correct and reasonably efficient solutions to programming problems Austin et al. (2021); Chen et al. (2021), among many other code-related skills Li et al. (2023). Shypula et al. (2023) have shown that CLMs can optimise slow-running code to achieve large runtime gains but at a substantial cost to correctness (down by up to $\sim$30%). Subsequent research has focused mostly on improving code correctness. From the data perspective, a common way of improving functional correctness is via distilled supervised fine-tuning Tunstall et al. (2023); Xu et al. (2023); Luo et al. (2023); Wei et al. (2023) on training data generated by large models such as GPT-4 Achiam et al. (2023). However, in many cases, due to legal, financial and/or privacy constraints, it is not feasible to rely on proprietary data. Furthermore, we seek to overcome the limitations of supervised fine-tuning (SFT), which only optimises for ‘positive’ examples and possesses no means of reducing the likelihood that undesirable (e.g. incorrect or slow) code is generated. Although such issues may be addressed via Reinforcement Learning (RL; Le et al. (2022); Wang et al. (2022); Gorinski et al. (2023)), they often come with added complexity and instability. Therefore, we opt for Direct Preference Optimisation Rafailov et al. (2024) as our preferred fine-tuning method due to its simplicity and widespread adoption.
We propose Code-Optimise, a lightweight framework that trains CLMs with our self-generated preference data for correctness (passed / failed) and efficiency (quick / slow), shown in Figure 1. Starting from a small collection of problems and unit tests, Code-Optimise bootstraps the pre-trained CLM to generate the required learning signals, thereby exposing the model to automatically annotated on-policy data. Additional robustness is also provided by dynamically selecting solutions during training to reduce overfitting. Our method consists of three steps: 1) Sampling; generate N solutions for each problem description, 2) Annotation; automatically label each solution for correctness and runtime, 3) Optimisation; fine-tune the CLM on self-generated preference data using several lightweight configurations. Our main contributions are:
  • We create and publish a novel code preference dataset (and recipe to extend it) that enables multi-objective optimisation (code correctness and runtime efficiency) of CLMs.
@@
  • We demonstrate that runtimes are reduced by up to 6% for MBPP and 3% for HumanEval over competitive baseline CLMs. Finally, the length of generated solutions is reduced by up to 23% for HumanEval and 48% for MBPP, thereby decreasing inference costs.
@@ 2 Code-Optimise
In Figure 1, we introduce Code-Optimise, a lightweight optimisation method for CLMs aimed at improving functional correctness and/or reducing the runtime of code. We assume that the pre-trained model has knowledge of accurate and efficient solutions but requires guidance to determine the best solution. This is also one of the main motivations of preference optimisation algorithms. Unlike training with distilled signals from larger models, new information is not learned. These approaches may be orthogonal. However, such an exploration is beyond the scope of this work.
@@ 2.1 Sampling
We assume access to $D_{seed}=\{x_{i},y_{i},ut_{i}\}_{i=1}^{N}$, a dataset of problem descriptions $x_{i}$ and the corresponding unit tests $ut_{i}$ that can be used for sampling and testing new solutions from the CLM, denoted $CLM_{base}$ henceforth. Since fine-tuning the model on the limited solutions $y_{i}$ would lead to overfitting, we leverage its extensive pre-training to generate a multitude of diverse solutions to obtain additional training data. We sample 100 solutions from $CLM_{base}$ for each problem description with multinomial sampling due to its lower computational cost. A temperature of $t=0.6$ is applied to achieve a balance between functional correctness and diversity, resulting in non-uniform runtimes.
@@ 2.3 Optimisation
In this step, the model is fine-tuned on $D_{train}$ to bias $CLM_{base}$ towards generating more functionally correct and runtime-efficient solutions. Although several methods for preference data optimisation exist (Yuan et al., 2023; Zhao et al., 2023; Liu et al., 2023; Azar et al., 2024; Ethayarajh et al., 2024; Hong et al., 2024), we opt for DPO due to its simplicity and wide adoption. We also benchmark SFT due to its widespread use in prior work.
@@ 3.4 Model Scaling
Figures 4, 5 and 6 show the evolution of functional correctness, runtimes and lengths of generated solutions as the number of trainable parameters increases (1B - 13B). Analysing $pass@1$ in Figure 4, we can see that larger DPO models achieve a more significant improvement over the baseline and SFT, particularly for in-domain problems. Somewhat surprisingly, functional correctness for HumanEval (out-of-domain) improves at a faster rate than MBPP (up to 7B parameters). In Figure 5, we observe that as the DPO models increase in size, their runtimes relative to the baseline remain consistent. The $DPO_{PvF}$ configuration tends to average somewhat slower runtimes as this setup only optimises for correctness thus sacrificing efficiency. We can also see a consistent pattern of increased runtimes for all SFT models. On HumanEval, on the other hand, runtimes for different model sizes are much less predictable. However, on average, our best configuration $DPO_{QvS}$ does show an improvement over the already competitive baseline CLMs. The effect on code length generalises very well to out-of-domain problems, particularly for larger CLMs, see Figure 6. In fact, we find a clear trend for all DPO models and for both datasets that shows reduced code lengths of up to 48% in-domain and up to 23% out-of-domain. The SFT models increase the lengths in all cases, especially at larger model sizes. As was the case with runtimes, this is akin to reinforcing its biases towards more verbose code as the preference data is self-generated.
@@ 3.5 Qualitative Analysis
Figure 8 shows several solutions to a typical programming problem taken from MBPP that gives a more tangible form to our results. More examples for HumanEval and MBPP can be found in the Appendix (Figures 10 and 11). Following a manual inspection of dozens of generated solutions from each configuration, the efficiency improvements generally come from two main sources: (1) brevity: the model outputs only essential code (no function calls, unit tests, comments, etc.), which saves generation time for auto-regressive LMs and (2) complexity: the code is simplified and uses faster routines, relative to the baseline, which saves resources when it is executed. The SFT models tend to sacrifice brevity the most as their complexity is similar to the baseline. Figures 8, 10, and 11 show several examples of this, e.g. adding function calls to the newly generated solution, possibly with import statements and/or expected inputs or outputs. [^1] Adding test cases to code is generally considered good programming practice. However, for the purpose of pure efficiency (our case), this can potentially add significant cost. This is in line with the observation from Figure 3 where the SFT models appear to be more verbose and biased towards longer outputs. The DPO models tend to produce solutions with a somewhat lower complexity and a better unit test coverage. Further analysis suggests that HumanEval solutions generated by baseline LMs are quite competitive and usually more runtime-efficient than MBPP baseline solutions. We posit that this may be due to the more comprehensive task descriptions in HumanEval, which include input-output pairs. Among DPO models, we do not observe a clear winner in qualitative analysis although $DPO_{QvS}$ is the best setup in terms of aggregate results.
3.6 Importance of Solution Selection
3.7 Which configuration to choose?
Any configuration that optimises for a single objective is expected to perform best in said objective. As such, when maximising functional correctness, DPOPvF is preferred. On the other hand, DPOQvS works best for runtime efficiency. With DPOAll, the model improves in functional correctness and runtime efficiency but is not optimal at both. Practitioners should choose the configuration that works best for their given code generation objective.
4 Related Work
Distilled Supervised Fine-Tuning
has been applied to code generation for the sole purpose of improving functional correctness. CLMs such as MagiCoder Wei et al. (2023) and WizardCoder Luo et al. (2023) query large proprietary models to provide the necessary training data Cui et al. (2024); Xu et al. (2023). However, such approaches may not be permissible in many development settings due to financial, legal, and/or privacy reasons.
Reinforcement Learning
Le et al. (2022); Wang et al. (2022); Gorinski et al. (2023) can overcome the shortcomings of supervised fine-tuning by propagating the negative rewards for dysfunctional code. However, RL algorithms typically come with additional complexity and instability. Recently, Rafailov et al. (2024) proposed Direct Preference Optimisation (DPO) as an alternative to Reinforcement Learning from Human Feedback when aligning language models to human preferences Tunstall et al. (2023). DPO serves as a form of offline RL that directly optimises on a given set of trajectories without requiring a separate reward model. Note that the aforementioned RL approaches only consider code correctness and not runtime as well.
Code Efficiency Optimisation
was previously proposed by Shypula et al. (2023) as a code editing task where slow-running code is altered to achieve a faster runtime. Optimisation was done on a newly curated dataset with synthetic test cases from various methods of prompting and fine-tuning. However, the greatly reduced runtimes came at a significant cost to code correctness. In many configurations, performance was reduced by up to 30%, with ‘smaller’ CLMs (7B, 13B) suffering a larger degradation. We hypothesise that this may be due to a) overfitting the single runtime objective (unlike our work where we aim to optimise for correctness and runtime) and b) removing failed programs from the dataset, thus leading to CLMs that struggle with the semantics of correct vs. incorrect code. We opt to not compare directly with this work as their method was specifically curated for code editing where a correct but inefficient program is assumed as input. On the contrary, we seek to produce solutions from scratch that are both correct and runtime efficient.
@@ 5 Conclusions
We have introduced Code-Optimise, a lightweight framework for improving code correctness and runtime via self-generated preference data (quick vs. slow and passed vs. failed solutions). Our experiments have shown several benefits: 1) functional correctness is significantly improved, particularly for smaller models, 2) dynamic solution selection during training provides an additional benefit by reducing overfitting, 3) runtime is reduced by up to 6% for MBPP and 3% for HumanEval over the baseline CLM, lowering the cost of code execution, 4) code length is significantly shortened, up to 48% for MBPP and 23% for HumanEval, which reduces inference cost and improves response time. We hope that our insights and novel dataset will stimulate further exciting research in this area.
@@ 6 Limitations
Timing the execution of short programs accurately is challenging and despite our best efforts, the runtime measurements could probably be improved with additional software engineering. This would also provide a cleaner and more stable learning signal for Code-Optimise. While our methodology is highly data-efficient, using only $\sim$200 open-source prompts for generating the training data, obtaining high-quality problems (free from proprietary/licensing issues) may yield better results. For code-related tasks that are amenable to our methodology, the improvements to runtime/inference may be investigated outside the scope of this paper. While we conducted all experiments using Python, we acknowledge that other programming languages should also be analysed in follow-up work.
@@ A.1 Dataset
The Mostly Basic Programming Problems introduced by MBPP consists of 974 crowd-sourced Python programming challenges. Each problem comprises a description, an example code solution and a few automated test cases. The dataset contains training, validation and test splits. We utilise the training and validation splits for optimisation, while the test split serves as the in-domain test data distribution.
@@
HumanEval comprises 164 Python programming challenges. The function signatures, docstrings, example solutions and several unit tests were handwritten for each problem. We leverage HumanEval as our out-of-domain test set as the descriptions in MBPP do not contain any unit tests and the writing style of HumanEval problems does not follow a consistent format. This helps us evaluate robustness to handwritten prompts.
@@ A.2 Training
We use the StarCoder StarCoder and CodeLlama CodeLlama families of models in our experiments. We opt for the pretrained (base) versions with sizes of 1B and 3B for StarCoder and 7B and 13B for CodeLlama, hosted on HuggingFace Wolf et al. (2020). During training, we fine-tune each model using a total of 30 epochs and select the best model based on the lowest validation loss. We use a learning rate of $5e^{-7}$ with a linear scheduler, a 10% warm-up, and a maximum sequence length of 2048 tokens.